# Measuring Chain of Thought Faithfulness by Unlearning Reasoning Steps

### v2

@@ 1 Introduction
Figure 1: An illustration of fur. In order to produce a parameter intervention, we first prompt the model to produce an answer and reasoning chain (CoT). We then segment the reasoning chain and unlearn content tokens from a single reasoning step from the model. The unlearned model is then prompted to produce an answer. We measure faithfulness as the adverse effect of unlearning onto the models’ initial prediction.
@@
Language models (LMs) can perform various tasks accurately and verbalize some reasoning via a so-called chain of thought (CoT) Kojima et al. (2022); Wei et al. (2022), even without specialized supervised training. CoT reasoning is emerging as a powerful technique for improving the performance of LMs in complex tasks OpenAI (2024); Guo et al. (2025). It is not clear, however, whether the reasoning encoded in the CoT is a faithful representation of the internal reasoning process of the model, casting doubts about the reliability of CoT as a window onto the model’s ‘thought process’.
@@
Various works set out to explore CoT faithfulness by perturbing tokens within the CoT and observing whether the contextual corruptions affect model prediction (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024; Madsen et al., 2024). This setup is inherently imprecise, as erasing steps from context does not remove knowledge from model parameters, and the model may still be able to reconstruct corrupted information when generating a prediction. Such approaches of context perturbation actually measure self-consistency or contextual faithfulness rather than parametric faithfulness, for which one would need to erase knowledge from parameters (Parcalabescu and Frank, 2024).
@@
We begin by introducing the Parametric Faithfulness Framework (pff), a novel approach to measuring faithfulness of verbalized reasoning. We define necessary components of instances of such a framework in two stages: (1) an intervention on the model parameters; and (2) evaluating parametric faithfulness. See components in Figure 1. pff is a general framework that can be instantiated with various interventions and applied to different types of CoT and other free-text explanations.
@@
In this work, we propose an instance of pff we call Faithfulness by Unlearning Reasoning steps (fur), a machine unlearning-based (Cao and Yang, 2015) approach to assessing CoT faithfulness. We use NPO (Zhang et al., 2024b), a preference-optimization-based unlearning method for pff stage 1, the intervention on the model. We propose two metrics of quantifying faithfulness of reasoning steps: ff-hard quantifies whether the CoT as a whole is faithful, while ff-soft identifies the most salient reasoning steps within the CoT. Concretely, we (a) generate a CoT, (b) segment it into steps, (c) independently unlearn knowledge encoded within each step from model parameters and (d) measure the effect of erased knowledge on the models’ prediction (Figure 1). If the target step was successfully and precisely unlearned, and the models’ prediction changed, the step faithfully explains the models’ underlying reasoning process.
@@
Through experimental evaluation on four LMs and five MCQA multi-hop reasoning datasets, we show we are able to perform valid interventions affect model predictions while not damaging the models’ general capabilities. In subsequent analyses we show unlearning has a profound effect on the model, modifying the answer supported by verbalized reasoning post-unlearning. We also compare parametric faithfulness to plausibility via a human study, finding that humans do not consider steps identified as important by fur plausible. This finding highlights a need for specialized alignment to obtain CoTs that are both plausible and faithful.
@@ 2 Background and Related Work
When CoT prompted, models exhibit better performance on complex multi-hop and arithmetic reasoning tasks (Zhou et al., 2023; Fu et al., 2023b; Sprague et al., 2024) compared to being prompted directly (no-CoT). Chains of thought can be used as additional context where models can store results of intermediate hops, but they also provide additional compute irrespective of content Pfau et al. (2024); Biran et al. (2024). Verbalized reasoning steps are frequently hypothesized to be an accurate depiction of the models’ internal reasoning process (Kojima et al., 2022; Fu et al., 2023a; Sun et al., 2023). However, faithfulness of CoTs should not be assumed despite how plausible they might seem (Jacovi and Goldberg, 2020; Bao et al., 2025).
@@
Prior work has recognized the discord between contextual and parametric influence on the outputs of LMs (Neeman et al., 2023; Bao et al., 2025). Prompting models with hypothetical or factually incorrect information causes them to change their otherwise consistently correct predictions (Kim et al., 2021, 2023; Simhi et al., 2024; Minder et al., 2024), highlighting their high sensitivity to context tokens and confounding any conclusions drawn from contextual perturbations applied to reasoning steps. The main issue with work investigating self-consistency is the possibility of the LM reconstructing information obfuscated by the contextual perturbation—despite the verbalized knowledge missing, this reasoning could still be retrieved from the latent space (Yang et al., 2024; Deng et al., 2024). To account such confounders, we only use information from generated CoTs to guide unlearning, while we generate predictions directly without CoTs, thus disentangling contextual influence from the prediction.
@@
Various tests and metrics for quantifying faithfulness of free-text explanations in LMs have previously been proposed (Lanham et al., 2023; Bentham et al., 2024; Atanasova et al., 2023; Siegel et al., 2024). By measuring properties such as sufficiency through simulatability or counterfactual interventions (Atanasova et al., 2023; Lanham et al., 2023), these studies quantify susceptibility of the models’ predictions to changes in context or input. Such approaches are valid only if there is no direct causal link between the input and prediction that bypasses the explanation, which is rare in LMs (Bao et al., 2025). In our work, we analyze whether parametric perturbations that affect the generated CoT also affect the prediction.. The closest to ours are the contemporaneous works of Yeo et al. (2024) who use activation patching to measure causal effect of corrupting certain hidden states, and Zaman and Srivastava (2025) who use knowledge editing to evaluate existing (un)faithfulness metrics.
@@
Stage 1: InterventionStage 2: EvaluationUnlearning (§4.1)NPO+KL (Eq. 9)on individual CoT stepsControls (§4.2)1. Efficacy (Eq. 2) 2. Specificity (Eq. 3) 3. General capabilities $\Delta$prediction (§6.2)Prediction differencebetween $\mathcal{M}$ and $\mathcal{M}^{*}{}$.$\text{ff}_{\text{hard}}$ and $\text{ff}_{\text{soft}}$4.3)$\Delta$reasoning (§6.3)CoT Differencebetween $\mathcal{M}$ and $\mathcal{M}^{*}{}$
@@
A line of work has analyzed the sensitivity of models to perturbations applied to reasoning steps (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024; Madsen et al., 2024, inter alia) under the guise of faithfulness. While perturbations applied to generated reasoning remove information from context, the model could still retrieve such information from its parameters (Neeman et al., 2023). Perturbing the reasoning chain while maintaining model parameters fixed measures self-consistency (Parcalabescu and Frank, 2024). Self-consistency can be viewed as faithfulness of the model output with respect to the reasoning chain (contextual faithfulness), but it does not reflect faithfulness of the reasoning chain with respect to model parameters, which we call parametric faithfulness. Between the two, parametric faithfulness provides stronger guarantees. Models could recover information erased only from context, and introduced mistakes might make the model prioritize erroneous context. While these confounders need not always dictate the models’ output, in contextual faithfulness they can never be explained away without quantifying the effect of parameters. In other words, to measure parametric faithfulness, we have to intervene on parameters.
@@
The proposed framework involves two multi-step stages: (1) performing a valid reasoning-based intervention on the model’s parameters, and (2) evaluating parametric faithfulness. We outline our framework in Figure 2.
@@
In the second stage, faithfulness is assessed with at least one of two evaluation protocols: (1) Instruct both $\mathcal{M}^{*}$ and $\mathcal{M}$ to directly give answers, then compute how often and how strongly their answers differ. (2) Instruct $\mathcal{M}^{*}$ and $\mathcal{M}$ to reason-then-answer, then compute how often they present different reasoning. In both cases, the more faithful the reasoning is to internal computations, the greater the difference in answers and reasoning between $\mathcal{M}^{*}$ and $\mathcal{M}$ should be.
@@ 4.1 Parameter Intervention
We erase knowledge contained in the verbalized reasoning steps using a preference-optimization unlearning method, NPO (Zhang et al., 2024b). Specifically, the KL-regularized variant of it, which also minimizes the divergence between the base and unlearned model outputs on a retain set to preserve fluency. [^2] We experimented with NPO+grad-diff, but results were always slightly worse than NPO+KL. We explored ROME and MEMIT (Meng et al., 2022, 2023), but they require a structured format, and do not perform well under paraphrases. We provide a detailed overview in Appendix A for readers unfamiliar with NPO+KL.
NPO+KL requires defining the forget set, $\mathcal{D}_{\text{FG}}$, and the retain set, $\mathcal{D}_{\text{RT}}$, which we construct as follows. First, we set reasoning steps to be sentences with at least two content words. For each step, we construct its $\mathcal{D}_{\text{FG}}$ of input-output pairs formed by taking, for each content word in the step, the prefix up to that word as input and the tokens of the content word as output. [^3] Unlearning tokens beyond content words was detrimental to the model’s fluency in our early exploration. NPO updates the model’s parameters to discourage it from predicting content words $\mathcal{D}_{\text{FG}}$ given prior context. We similarly construct $\mathcal{D}_{\text{RT}}$ for a given step as content words from four randomly selected CoT steps from other instances. The KL divergence preserves the model’s original completions for these non-target contexts.
We unlearn each reasoning step individually, for a total of $5$ iterations, and refer to the model obtained after unlearning the $i$-th reasoning step alone as $\mathcal{M}^{(i)^{*}}$. One unlearning iteration refers to a pass over $\mathcal{D}_{\text{FG}}$. We only update the second FF2 matrix of the Transformer MLPs, as this layer was found to act as a memory store (Geva et al., 2021b; Meng et al., 2022) and model editing methods frequently target it to update information (Meng et al., 2022, 2023; Hong et al., 2024). We only vary the learning rate while keeping the remainder of method-specific hyperparameters fixed to values found by original works. We report them in Appendix E.
@@ 4.2 Controls
In order to measure whether unlearning affects general model capabilities, we compare the performance on MMLU (Hendrycks et al., 2021) before and after unlearning. Due to prohibitive costs of evaluating few-shot MMLU for each instance and unlearned CoT step, we (1) opt for zero-shot evaluation as the instruction-tuned models report good performance in this setup, and (2) report full MMLU scores on a randomly selected sample of $10$ CoTs after unlearning each ($\approx 50$) CoT step from models.
@@
Note that we do not aim for efficacy to reach $1$, as that would imply that the unlearned step has probability 00 (Eq. 2), which in turn would likely adversely affect the fluency of the model. Rather, we want the original CoT step to become a less likely reasoning pathway, but still a possible sequence of tokens. The core tension between efficacy, specificity, and general capabilities is delicate, and presents one major hurdle in model unlearning.
@@ 4.3 Faithfulness Measurements
Perfectly determining whether a reasoning chain constitutes a faithful explanation is difficult. Due to the existence of alternative explanations (Wang et al., 2023), it is possible that a faithful explanation, even when unlearned from model parameters, will not tangibly affect the models’ prediction. Therefore, we do not expect $\text{ff}_{\text{hard}}$ to have perfect recall. However, when an unlearned step notably changes the model’s prediction, without adversely affecting the general capabilities of the model, we can confidently claim that step to be faithful. For the remaining $100-\text{ff}$ instances, there are three possibilities: (1) fur failed to uncover and unlearn the true reasoning path, (2) the model used multiple valid reasoning paths, and unlearning one did not significantly affect its prediction, or (3) the model was genuinely unfaithful in its explanation. In this sense, ff represents a lower bound on the model’s true faithfulness — it is the rate at which we can successfully uncover faithful reasoning (assuming that the flip happened due to a valid intervention).
@@ 5 Experimental Setup
We employ five diverse multi-hop datasets: OpenbookQA (Book; Mihaylov et al., 2018), ARC-Challenge (Arc-ch; Clark et al., 2018), StrategyQA (SQA; Geva et al., 2021a), TruthfulQA (TQA; Lin et al., 2022) and the Sports understanding subtask of BigBench-Hard Srivastava et al. (2023). These datasets span a variety of domains, necessitating knowledge of science, sports, geography, health, law, finance and logic. We choose MCQA as the target task as it simplifies analysis of how the models’ predictive distribution shifts after unlearning due to availability of alternative answers. To retain comparable sizes, and due to expensive runtime of unlearning each CoT step, we select a subset of $250$ instances from the test split of each dataset to balance the question sources. [^6] For SQA, we use instances from the validation split due to the availability of labels. Sports has a total of $248$ instances. Details of datasets and models are in Appendix C.
@@
We use a two-step prompting approach Bowman et al. (2022); Lanham et al. (2023), where the model is first prompted to generate the CoT based on the question and answer options, and subsequently prompted to complete the answer letter based on the question, answer choices, and the CoT. We use greedy decoding when generating, producing a single CoT for each instance. For the prompts used, see Appendix D.
@@
Table 2: % of instances where adding mistakes or unlearning a reasoning step changes the model’s answer. Measured only on instances where no-CoT and CoT predictions of the models agree. Scores over $1\%$ better in bold.

### v3

@@ 1 Introduction
Figure 1: An illustration of pff and fur. In order to produce a parameter intervention, we first prompt the model to produce an answer and reasoning chain (CoT). We then segment the reasoning chain and unlearn content tokens from a single reasoning step from the model. The unlearned model is then prompted to produce an answer. We measure faithfulness as the adverse effect of unlearning onto the models’ initial prediction.
@@
Language models (LMs) can perform various tasks accurately and verbalize some reasoning via a so-called chain of thought (CoT) Kojima et al. (2022); Wei et al. (2022), even without specialized supervised training. CoT reasoning is emerging as a powerful technique for improving the performance of LMs in complex tasks OpenAI (2024); DeepSeek-AI et al. (2025). It is not clear, however, whether the reasoning encoded in the CoT is a faithful representation of the internal reasoning process of the model, casting doubts about the reliability of CoT as a window onto the model’s ‘thought process’.
@@
Various works set out to explore CoT faithfulness by perturbing tokens within the CoT and observing whether the contextual corruptions affect model prediction (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024b; Madsen et al., 2024). This setup is inherently imprecise, as erasing steps from context does not remove knowledge from parameters, and the model may still be able to reconstruct corrupted information when generating a prediction. Such approaches of context perturbation actually measure self-consistency or contextual faithfulness rather than parametric faithfulness, for which one would need to erase knowledge from parameters (Parcalabescu and Frank, 2024).
@@
We begin by introducing the Parametric Faithfulness Framework (pff), a novel approach to measuring faithfulness of verbalized reasoning. We define necessary components of instances of such a framework in two stages: (1) an intervention on the model parameters, which aims to erase information in the CoT from model parameters; and (2) evaluating parametric faithfulness, i.e. whether the intervention affected the models’ prediction. See components in Figure˜1. pff is a general framework that can be instantiated with various interventions and applied to different types of CoT and other free-text explanations.
@@
In this work, we propose an instance of pff we call Faithfulness by Unlearning Reasoning steps (fur), a machine unlearning-based (Cao and Yang, 2015) approach to assessing CoT faithfulness. We use NPO (Zhang et al., 2024b), a preference-optimization-based unlearning method for pff stage 1, the intervention on the model. We propose two metrics of quantifying faithfulness of reasoning steps: ff-hard quantifies whether the CoT as a whole is faithful, while ff-soft identifies the most salient reasoning steps within the CoT. Concretely, we (a) generate a CoT, (b) segment it into steps, (c) independently unlearn knowledge encoded within each step from model parameters and (d) measure the effect of erased knowledge on the models’ prediction (Figure˜1). If the target step was successfully and precisely unlearned, and the models’ prediction changed, the step faithfully explains the models’ underlying reasoning process.
@@
Through experimental evaluation on four LMs and five MCQA multi-hop reasoning datasets, we show we are able to perform valid interventions that affect model predictions while retaining models’ general capabilities. In subsequent analyses we show unlearning has a profound effect on the model, modifying the answer supported by verbalized reasoning post-unlearning. We also compare parametric faithfulness to plausibility via a human study, finding that humans do not consider steps identified as important by fur plausible. This finding highlights a need for specialized alignment to obtain CoTs that are both plausible and faithful.
@@ 2 Background and Related Work
When CoT prompted, models exhibit better performance on complex multi-hop and arithmetic reasoning tasks (Zhou et al., 2023; Fu et al., 2023b; Sprague et al., 2025) compared to being prompted directly (no-CoT). Chains of thought can be used as additional context where models can store results of intermediate hops, but they also provide additional compute irrespective of content Pfau et al. (2024); Biran et al. (2024). Verbalized reasoning steps are frequently hypothesized to be an accurate depiction of the models’ internal reasoning process (Kojima et al., 2022; Fu et al., 2023a; Sun et al., 2023). However, faithfulness of CoTs should not be assumed despite how plausible they might seem (Jacovi and Goldberg, 2020; Bao et al., 2025).
@@
Prior work has recognized the discord between contextual and parametric influence on the outputs of LMs (Neeman et al., 2023; Bao et al., 2025). Prompting models with hypothetical or factually incorrect information causes them to change their otherwise consistently correct predictions (Kim et al., 2021, 2023; Simhi et al., 2024; Minder et al., 2025), highlighting their high sensitivity to context tokens and confounding any conclusions drawn from contextual perturbations applied to reasoning steps. The main issue with work investigating self-consistency is the possibility of the LM reconstructing information obfuscated by the contextual perturbation—despite the verbalized knowledge missing, this reasoning could still be retrieved from the latent space (Yang et al., 2024; Deng et al., 2024). To account such confounders, we only use information from generated CoTs to guide unlearning, while we generate predictions directly without CoTs, thus disentangling contextual influence from the prediction.
@@
Various tests and metrics for quantifying faithfulness of free-text explanations in LMs have previously been proposed (Lanham et al., 2023; Bentham et al., 2024; Atanasova et al., 2023; Siegel et al., 2024). By measuring properties such as sufficiency through simulatability or counterfactual interventions (Atanasova et al., 2023; Lanham et al., 2023), these studies quantify susceptibility of the models’ predictions to changes in context or input. Such approaches are valid only if there is no direct causal link between the input and prediction that bypasses the explanation, which is rare in LMs (Bao et al., 2025). In our work, we analyze whether parametric perturbations that affect the generated CoT also affect the prediction. The closest to ours are the contemporaneous works of Yeo et al. (2024) who use activation patching to measure causal effect of corrupting certain hidden states, and Zaman and Srivastava (2025) who use knowledge editing to evaluate existing (un)faithfulness metrics.
@@
[…]
@@
A line of work has analyzed the sensitivity of models to perturbations applied to reasoning steps (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024b; Madsen et al., 2024, inter alia) under the guise of faithfulness. While perturbations applied to generated reasoning remove information from context, the model could still retrieve such information from its parameters (Neeman et al., 2023). Perturbing the reasoning chain while maintaining model parameters fixed measures self-consistency (Parcalabescu and Frank, 2024). Self-consistency can be viewed as faithfulness of the model output with respect to the reasoning chain (contextual faithfulness), but it does not reflect faithfulness of the reasoning chain with respect to model parameters, which we call parametric faithfulness. Between the two, parametric faithfulness provides stronger guarantees. Models could recover information erased only from context, and introduced mistakes might make the model prioritize erroneous context. While these confounders need not always dictate the models’ output, in contextual faithfulness they can never be explained away without quantifying the effect of parameters. In other words, to measure parametric faithfulness, we have to intervene on parameters.
@@
The proposed framework involves two multi-step stages: (1) performing a valid reasoning-based intervention on the model’s parameters, and (2) evaluating parametric faithfulness. We outline our framework in Figure˜2.
@@
In the second stage, faithfulness is assessed with at least one of two evaluation protocols: (1) Instruct both $\mathcal{M}^{*}$ and $\mathcal{M}$ to directly give answers, then compute how often and how strongly their answers differ; (2) Instruct $\mathcal{M}^{*}$ and $\mathcal{M}$ to reason-then-answer, then compute how often they present different reasoning. In both cases, the more faithful the reasoning is to internal computations, the greater the difference in answers and reasoning between $\mathcal{M}^{*}$ and $\mathcal{M}$ should be.
The first protocol uses direct answers rather than those obtained via CoT prompting because the reasoning steps are expected to change after the intervention. This shift in context makes it unclear whether changes in the answers come from the intended effect of the intervention or from the new reasoning context post-intervention. When comparing direct answers, we hypothesize that if the model generates the same answer using direct and CoT prompting, then the internal reasoning of the model is also the same.
@@ 4.1 Parameter Intervention
We erase knowledge contained in the verbalized reasoning steps using a preference-optimization unlearning method, NPO (Zhang et al., 2024b). Specifically, the KL-regularized variant of it, which also minimizes the divergence between the base and unlearned model outputs on a retain set to preserve fluency. [^2] We experimented with NPO+grad-diff, but results were always slightly worse than NPO+KL. We explored ROME and MEMIT (Meng et al., 2022, 2023), but they require a structured format, and do not perform well under paraphrases. We provide a detailed overview in Appendix˜A for readers unfamiliar with NPO+KL.
NPO+KL requires defining the forget set, $\mathcal{D}_{\text{FG}}$, and the retain set, $\mathcal{D}_{\text{RT}}$, which we construct as follows. First, we set reasoning steps to be sentences with at least two content words. For each step, we construct its $\mathcal{D}_{\text{FG}}$ of input-output pairs formed by taking, for each content word in the step, the prefix up to that word as input and the tokens of the content word as output. [^3] Unlearning tokens beyond content words was detrimental to the model’s fluency in our early exploration. NPO updates the model’s parameters to discourage it from predicting content words $\mathcal{D}_{\text{FG}}$ given prior context. We similarly construct $\mathcal{D}_{\text{RT}}$ for a given step as content words from four randomly selected CoT steps from other instances. Concretely, we sample $4$ other instances from the same dataset and randomly select a CoT step with at least two content words from each instance. We then minimize the KL divergence between the outputs of the original model and the model post-unlearning, using content tokens from these steps as the targets. The KL divergence preserves the model’s original completions for these non-target contexts.
We unlearn each reasoning step individually, for a total of $5$ iterations, and refer to the model obtained after unlearning the $i$-th reasoning step alone as $\mathcal{M}^{(i)^{*}}$. One unlearning iteration refers to a pass over $\mathcal{D}_{\text{FG}}$. We only update the second FF2 matrix of the Transformer MLPs, as this layer was found to act as a memory store (Geva et al., 2021b; Meng et al., 2022) and model editing methods frequently target it to update information (Meng et al., 2022, 2023; Hong et al., 2024). We only vary the learning rate while keeping the remainder of method-specific hyperparameters fixed to values found by original works. We report them in Appendix˜E.
@@ 4.2 Controls
In order to measure whether unlearning affects general model capabilities, we compare the performance on MMLU (Hendrycks et al., 2021) before and after unlearning. Due to prohibitive costs of evaluating few-shot MMLU for each instance and unlearned CoT step, we (1) opt for zero-shot evaluation as the instruction-tuned models report good performance in this setup, and (2) MMLU score of the model after unlearning each step from $10$ randomly selected CoTs ($\approx 50$ unlearning steps).
@@
Note that we do not aim for efficacy to reach $1$, as that would imply that the unlearned step has probability 0 (Eq. 2), which in turn would likely adversely affect the fluency of the model. Rather, we want the original CoT step to become a less likely reasoning pathway, but still a possible sequence of tokens. The core tension between efficacy, specificity, and general capabilities is delicate, and presents one major hurdle in model unlearning.
@@ 4.3 Faithfulness Measurements
Perfectly determining whether a reasoning chain constitutes a faithful explanation is difficult. Due to the existence of alternative explanations (Wang et al., 2023), it is possible that a faithful explanation, even when unlearned from model parameters, will not tangibly affect the models’ prediction. Therefore, we do not expect ff-hard to have perfect recall. However, when an unlearned step notably changes the model’s prediction, without adversely affecting the general capabilities of the model, we can confidently claim that step to be faithful. For the remaining $100$-$f$-hard instances, there are three possibilities: (1) fur failed to uncover and unlearn the true reasoning path, (2) the model used multiple valid reasoning paths, and unlearning one did not significantly affect its prediction, or (3) the model was genuinely unfaithful in its explanation. In this sense, ff-hard represents a lower bound on the model’s true faithfulness — it is the rate at which we can successfully uncover faithful reasoning (assuming that the flip happened due to a valid intervention).
@@ 5 Experimental Setup
We employ five diverse multi-hop datasets: OpenbookQA (Book; Mihaylov et al., 2018), ARC-Challenge (Arc-ch; Clark et al., 2018), StrategyQA (SQA; Geva et al., 2021a), TruthfulQA (TQA; Lin et al., 2022) and the Sports understanding subtask of BigBench-Hard Srivastava et al. (2023). These datasets span a variety of domains, necessitating knowledge of science, sports, geography, health, law, finance and logic. We choose MCQA as the target task as it simplifies analysis of how the models’ predictive distribution shifts after unlearning due to availability of alternative answers. To retain comparable sizes, and due to expensive runtime of unlearning each CoT step, we select a subset of $250$ instances from the test split of each dataset to balance the question sources. [^6] For SQA, we use instances from the validation split due to the availability of labels. Sports has a total of $248$ instances. Details of datasets and models are in Appendix˜C.
@@
We use a two-step prompting approach Bowman et al. (2022); Lanham et al. (2023), where the model is first prompted to generate the CoT based on the question and answer options, and subsequently prompted to complete the answer letter based on the question, answer choices, and the CoT. We use greedy decoding when generating, producing a single CoT for each instance. For the prompts used, see Appendix˜D.
@@
Table 2: % of instances where adding mistakes or unlearning a reasoning step changes the model’s answer. Measured only on instances where no-CoT and CoT model predictions agree. Scores over $1\%$ higher in bold.